OpenStudy (anonymous):

Draw the set S of complex numbers such that \[|z-2i|<2,,, \Im (z)=1\] What is the argument when \[z \in S\]

OpenStudy (anonymous):

\[|a-i|<2\]

OpenStudy (anonymous):

But what does that say on a graph?

OpenStudy (anonymous):

-2<a-i<2

OpenStudy (experimentx):

says it lines in the region inside a circle with radius 2 and center 0,2

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i= |a-i|%3C2

OpenStudy (anonymous):

It's no circle?

OpenStudy (anonymous):

Is it because the Im(z) is constant at 1 and -2< Re(z) < 2, so it just varies in the Re(z) plane?

OpenStudy (experimentx):

so you want z such that it's imaginary part is constant?

OpenStudy (anonymous):

No i think I'm a bit wrong in my thoughts

OpenStudy (experimentx):

put z = x + i in that equation |x + i - 2i| < 2 x^2 + 1^2 < 4 x^2 < 3 |x| < sqrt(3) -3 < x < 3

OpenStudy (anonymous):

Is it not Re betweem -2 and 2 ?

OpenStudy (experimentx):

sorry ... sqrt(3) .. lol

OpenStudy (anonymous):

Ah, okey :) So how do I get the argument from that?

OpenStudy (anonymous):

Just to try -sqrt3-i and sqrt3-i ?

OpenStudy (anonymous):

But they are not included in the interval so instead should i take some random x value inside the interval?

OpenStudy (experimentx):

yeah you can get that ... |x| must be sqrt(3)'s ...just put it up and see what you get.

OpenStudy (anonymous):

So i should use -i as Im, right?

OpenStudy (anonymous):

If so the argument is \[\frac{ 11\pi }{ 6 }\]

OpenStudy (experimentx):

put z = x + i in that equation |x + i - 2i| < 2 x^2 + 1^2 < 4 x^2 < 3 |x| < sqrt(3) -sqrt(3) < x < sqrt(3) put z = x + 11pi/6 i <--- in above

OpenStudy (anonymous):

\[z=x+\frac{ 11\pi }{6 }i, \] where \[-\sqrt{3}<x<\sqrt{3}\]

OpenStudy (anonymous):

Is that what you meant?

OpenStudy (experimentx):

no ... find the value of x in similar fashion

OpenStudy (anonymous):

Sorry but i have to go for now. But thank you for your help and patience, will look back at this later.